OpenStudy (anonymous):

A tugboat goes 180 miles upstream in 15 hours. The return trip downstream takes 10 hours. find the speed of the tugboat without a current and the speed of the current.

OpenStudy (anonymous):

A tugboat goes 180 miles upstream in 15 hours. The return trip downstream takes 10 hours. Find the speed of the tugboat wihtout a current and the speed of the current. -------------------------- Upstream DATA: distance = 180 miles ; time = 15 hrs ; rate = 180/15 = 12 mph ----- Downstream DATA: distance = 180 miles ; time = 10 hrs ; rate = 180/10 = 18 mph ------ Equation: Upstream: b - c = 12 mph Downstrm: b + c = 18 mph

OpenStudy (anonymous):

thank you so much now i know how to work one like this if i have it again

OpenStudy (anonymous):

welcome make me best response